OpenStudy (anonymous):

What is the sum of 1/2 inch and 3/8 inch?

OpenStudy (anonymous):

1/2 + 3/8 = 4/8 + 3/8 = 7/8

OpenStudy (anonymous):

so 7/8

OpenStudy (anonymous):

covert 1/4 to 8ths

OpenStudy (anonymous):

as vikran did

OpenStudy (anonymous):

I am

OpenStudy (anonymous):

How do I convert though?

OpenStudy (anonymous):

I forgot how to!

OpenStudy (anonymous):

ok convert 1/2 to 8ths (not 1/4 - sorry) to get 8 on bottom we need to multiply the 2 by 4 but in order to make the fraction the same value we also have to multiply top by 4 as well s0 1 * 4 4 -- -- = -- 2 * 4 8

OpenStudy (anonymous):

peacelover - look at this plz!!!!

OpenStudy (anonymous):

so is it 4/8?

OpenStudy (anonymous):

is it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1/2 = 4/8

OpenStudy (anonymous):

How much longer than 7/8 inch is 3/4?

OpenStudy (anonymous):

dude which standard you are in peacelover ?

OpenStudy (anonymous):

convert 3/4 to 8ths and subtract fro 7/8

OpenStudy (anonymous):

how do I do that though?

OpenStudy (anonymous):

i've shown you how to convert 1/2 - now apply it to 3/4

OpenStudy (anonymous):

omg I am only in 6th grade go easy on me

OpenStudy (anonymous):

what do u multiply 4 by to get 8?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

right - so its 2 * 4 t0 get the 8 on bottom of fraction so we multiplied the bottom (thatis the 4 of the 3/4 to get 8) now we need to multiply the top part of fraction ( the 3) by 2 summarising: 3 * 2 = 6 -- -- -- 4 * 2 = 8 3/4 becomes 6 /8

OpenStudy (anonymous):

now we can subtract 7/8 - 6/8 = ( 7-6 )/8 = 1/8 its important to remember that you can only add or subtract fractions when the denominator ( the bottom of the fraction) is the same in both fractions

OpenStudy (anonymous):

so is it 1/8?

OpenStudy (anonymous):

hope this helps

OpenStudy (anonymous):

yes - the answer is 1/8

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw - try a few yourself - practice makes perfect

OpenStudy (anonymous):

thank god he understood... keep it up kid :)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

I'm a girl

OpenStudy (anonymous):

anyways you are a kid

OpenStudy (anonymous):

yeah in 6th grade!